# Synergy: Towards On-Body AI via Tiny AI Accelerator Collaboration on Wearables

### v2

@@ 5. Prototype Implementation
6. Evaluation
6.1. Experimental Setup
6.1.1. Workload
To demonstrate the effectiveness of Synergy orchestrating and executing on-body AI apps, we design four workloads with eight pipelines and eight different models (see Table 1). For the workload design, we considered two scenarios: concurrent apps, each using a different model (type, size) for different requirements (Workload 1 and 2) and a single app with a large model for enhanced capability (Workload 3 and 4). For the device setup, we used four MAX78000 devices, representing four smart wearables; an earbud, glasses, a watch, and a ring. We then assign source and target devices and a model based on the app scenarios. Figure 14 shows example execution plans of Workload 2 and 4. Pipeline 4 (keyword spotting), captures audio from an earbud ($d_{1}$), runs KWS (KWS20, [n. d.]), and sends results to a ring ($d_{4}$). Similarly, in Workload 4, pipeline 8 (object detector) captures images on glasses ($d_{2}$), runs MobileNetV2 (Sandler et al., 2018) across devices, and sends results to a ring ($d_{4}$). For each pipeline, various execution plans are generated by leveraging the four AI accelerators, including splitting models at different layers and assigning chunks to different devices.
Table 1. Workload and model description used in this study. Sizes are all in bytes.
@@ 6.1.1. Workload
2 ResSimpleNet(He et al., 2016) 381792 32$\times$32$\times$3 11217
3 UNet(Ronneberger et al., 2015) 279084 48$\times$48$\times$48 74547
2 4 KWS(KWS20, [n. d.]) 169472 128$\times$128$\times$1 7976
5 SimpleNet(Hasanpour et al., 2016) 166448 32$\times$32$\times$3 9237
6 WideNet(Hasanpour et al., 2016) 313700 32$\times$32$\times$3 10091
3 7 EfficientNetV2(Tan and Le, 2021) 627220 32$\times$32$\times$3 66468
4 8 MobileNetV2(Sandler et al., 2018) 821164 32$\times$32$\times$3 296318
Figure 14. Example execution plans (workload 2, 4).
Figure 15. Throughput, latency, and power consumption result for four different workloads.
6.1.2. Baselines
@@
IndModel: Inspired by state-of-the-art model partitioning methods (Kang et al., 2017; Hu et al., 2019; Zhang et al., 2020; Laskaridis et al., 2020; Jeong et al., 2018), IndModel selects the best-split execution plan based on metric estimations, but without a holistic view. IndModel independently selects the optimal execution plan for each pipeline and forms the final holistic collaboration plan by aggregating the selected plans.
@@
IndBest: IndBest advances IndModel by incorporating source and target devices alongside model splitting. Each pipeline independently selects the execution plan expected to yield the highest throughput by considering end-to-end latency, from sensing to output delivery. However, it does not account for the resource usage of other pipelines.
6.1.3. Performance Metrics
We consider three performance metrics: throughput, latency, and energy. Throughput is defined as the total number of model inferences per second, while latency is the time taken to execute an end-to-end holistic collaboration plan, respectively. For energy, we report the average power consumption for the execution of a holistic collaboration plan. We measured power consumption with Monsoon power monitor (Monsoon High Voltage Power Monitor, [n. d.]).
6.2. Overall Performance
Figure 15 shows the result of different methods for the four workloads. In terms of throughput, Synergy consistently outperforms all baselines owing to its holistic decision-making process and adaptive task parallelization. Synergy on average shows 23.0$\times$ higher throughput than the baselines. In Workload 1 and 2 where multiple pipelines run concurrently, IndModel (state-of-the-art model partitioning methods) results in OOR situations. In Workload 3 and 4, Synergy continues to exhibit 1.8$\times$ and 2.2$\times$ higher throughput than the second best (IndBest) respectively. The results indicate two key points. First, Synergy boosts throughput without OOR failures when multiple pipelines exist (Workload 1 and 2). Second, Synergy supports large models through the collaboration of distributed AI accelerators (Workload 3 and 4).
@@ 6.2. Overall Performance
Table 2. Ablation study of Synergy.
JRC STT PSR ATP TPUT (Wrkld 1) TPUT (Wrkld 2)
OOR OOR
0.06 2.30
0.92 15.28
2.72 15.28
4.20 29.67
6.2.1. Ablation Study
We perform an ablative study to understand the effectiveness of each technical component of Synergy. Table 2 shows the result. JRC refers to the joint resource consideration among pipelines, and STT refers to device mapping accounting for source and target tasks (§4.3). PSR means progressive search space reduction (§4.4), and ATP means adaptive task parallelization (§4.6). If none of these are applied, it is the same as the state-of-the-art approach (IndModel). If all are checked, it becomes Synergy. Overall, the throughput improves as each of our components is added. Improving IndModel by addressing resource conflicts among pipelines to avoid OOR is the same as JointModel in our baselines. Note that merely resolving the resource conflicts is still far from Synergy. Incorporating other technical components, STT, PSR, and ATP shows 6.9, 7.6, and 14.3$\times$ higher throughput compared to JRC-only, respectively.
6.3. In-Depth Analysis
6.3.1. Runtime Environment Changes
@@
Figure 16. Impact of runtime environment changes: (a) number of devices and (b) number of pipelines.
Number of devices: In this experiment, we explore Synergy’s scalability to the change of the number of MAX78000 devices. We increase the number of devices from two to five while running the same set of four pipelines with ConvNet5, KWS, SimpleNet, and ResSimpleNet. Figure 16 (a) shows the result. Interestingly, Synergy significantly outperforms the baselines as the number of devices increases. This is due to Synergy’s strategic consideration of accelerator assignments, effectively minimizing communication overhead between source and target devices. Conversely, except for IndBest, the throughput gains for all other baselines are not notable, even with additional devices. Another interesting observation is that using more devices does not always lead to higher throughput. In the case of Synergy, the throughput saturates once the number of devices reaches 4. This is because, beyond optimal distribution, further splitting of models fails to contribute to additional throughput gains.
@@
6.3.2. Composing Heterogeneous Accelerators
Figure 17. Effect of accelerator composition.
To examine Synergy’s effectiveness in heterogeneous accelerator resources, we conducted an experiment where one of four MAX78000 devices was substituted with more resource-capable MAX78002. The workload is comprised of three pipelines: ConvNet5, UNet, and EfficientNet. Figure 17 shows the results in two different setups. The inclusion of a higher-resource device generally led to an improvement in throughput. With four MAX78000s, Synergy achieves a total throughput of 0.93, outperforming the second best (JointModel) by a factor of 16.4. This throughput further increases to 3.33 when incorporating one MAX78002. PriMinDev, which allocates all models exclusively to the single MAX78002, results in a significantly low throughput of 0.06. This discrepancy highlights the importance of consideration for the communication overhead associated with source and target devices.
Figure 18. Effect of source and target mappings.
6.3.3. Source and Target Mapping
We assessed source and target mapping effects in three scenarios: (i) any device can be source or target (Any), (ii) source/target devices are evenly allocated (Distributed), and (iii) the same device is used as both source and target across pipelines (Overlapped). The Distributed setup matches Workload 1 from §6.2, while the Overlapped and Any scenarios only differ in their source and target device mappings. As shown in Figure 18, the Overlapped scenario has the lowest throughput due to communication bottlenecks, while the Any scenario has the highest throughput by distributing communication costs. The result indicates Synergy’s robustness in execution planning, considering both source and target factors.
Table 3. Comparison among different objectives.
@@ 6.3.4. Different Objectives
While Synergy defaults to maximizing system-wide throughput (TPUT-max), it can also minimize latency (Latency-min) or power consumption (Power-min). Table 3 demonstrates Synergy’s ability to achieve these different objectives effectively. Each objective excels in its metric, though identical latencies do not ensure the same holistic collaboration plan. Notably, maximizing throughput provides a balanced performance with low latency and near-minimal power use. For example, in workload 1, TPUT-max achieves 22.1$\times$ higher throughput with only 1.2$\times$ more power consumption compared to Power-min. This is because maximizing throughput reduces latency and communication overhead, the primary power consumption source.
7. Related work
@@
Model partitioning: While there are very few attempts for partitioning AI models over multiple MCU-equipped devices, there have been active research efforts for layer-wise model partitioning over resource-limited embedded and mobile devices. In common, these methods (Kang et al., 2017; Hu et al., 2019; Zhang et al., 2020; Laskaridis et al., 2020; Jeong et al., 2018; Wang et al., 2021; Banitalebi-Dehkordi et al., 2021) allocates few initial layers of a DNN on mobile/embedded device and the latter in edge or cloud server. Intermediate output from the initial layer execution is transmitted to powerful resources such as the cloud or nearby edge device where the subsequent part of the model is executed. They adapt the splitting layer depending on network status and server load. While Synergy also adopts vertical partitioning for distributed inferences, existing model partitioning studies lack the (i) end-to-end perspective of the on-body AI app execution (sensor and interface operations as well as model execution) and (ii) consideration of concurrent execution of multi-tenant models. Our experiments demonstrate that Synergy outperforms these studies by solving the challenges.
@@
DNN workload distribution: An application pipeline often consists of a (conditional) sequence of multiple DNN models. Several distributed systems have been proposed to balance model workloads across distributed devices (Zeng et al., 2020b; Jang et al., 2021; Dong et al., 2023). Their key difference with model partitioning methods is to treat a model as a primitive execution unit and focus more on scheduling model execution over distributed devices (without model splitting). For example, a face classification task can be executed on a different device following object detection on the initial device, thereby leveraging distributed resources from multiple devices (Zeng et al., 2020b; Jang et al., 2021). Synergy shares the same high-level objective of distributing multiple model workloads into multiple devices. However, our target environment is on-body devices with AI accelerators, which brings resource challenges. To overcome the limited resources of these devices, we adopt model partitioning on top of workload distribution and devise a tailored solution for tiny AI accelerators.
@@
Model serving systems: Several platforms, such as TensorFlow serving (TensorFlow Serving, [n. d.]), Sagemaker (SageMaker, [n. d.]), and Azure ML (ML as a Service, [n. d.]), have been proposed to facilitate model inference serving by offering containerized environments for model execution in diverse devices. Research platforms, such as Velox (Crankshaw et al., 2014) and Clipper (Crankshaw et al., 2017), focus on low-latency prediction serving, together with optimizing cloud server performance. However, in wearable computing, the main challenge lies in dynamically composing distributed tiny AI accelerators. Our device-agnostic programming interface addresses this by mapping software logic to physical resources and optimizing performance by considering pipeline interdependencies and resource usage.
8. Conclusion
@@
Future work aims to enhance collaboration among AI accelerators by expanding from layer-wise model splitting to include channel-wise splitting (Mao et al., 2017a, b; Zeng et al., 2020a; Zhao et al., 2018; Hou et al., 2022; Hu and Li, 2022), enabling Synergy to handle larger inputs and optimize performance through a combination of both techniques (Pang et al., 2023). Additionally, we plan to extend Synergy to support more complex model architectures, such as sensor fusion and conditional chaining of multiple models, to meet real-world application demands. Lastly, we aim to enrich the device-agnostic programming interface by adding features like detailed taxonomy for sensor and target device mapping and supporting customized functions, thereby providing developers with greater flexibility in designing on-body AI apps.

### v3

@@ V Prototype Implementation
VI Evaluation
VI-A Experimental Setup
VI-A1 Workload
To demonstrate the effectiveness of Synergy orchestrating and executing on-body AI apps, we design four workloads with eight pipelines and eight different models (see Table I). For the workload design, we considered two scenarios: concurrent apps, each using a different model (type, size) for different requirements (Workload 1 and 2) and a single app with a large model for enhanced capability (Workload 3 and 4). For the device setup, we used four MAX78000 devices, representing four smart wearables; an earbud, glasses, a watch, and a ring. We then assign source and target devices and a model based on the app scenarios. Figure 14 shows example execution plans of Workload 2 and 4. Pipeline 4 (keyword spotting), captures audio from an earbud ($d_{1}$), runs KWS [35], and sends results to a ring ($d_{4}$). Similarly, in Workload 4, pipeline 8 (object detector) captures images on glasses ($d_{2}$), runs MobileNetV2 [36] across devices, and sends results to a ring ($d_{4}$). For each pipeline, various execution plans are generated by leveraging the four AI accelerators, including splitting models at different layers and assigning chunks to different devices.
TABLE I: Workload and model description used in this study. Sizes are all in bytes.
@@ VI-A1 Workload
2 ResSimpleNet[37] 381792 32$\times$32$\times$3 11217
3 UNet[31] 279084 48$\times$48$\times$48 74547
2 4 KWS[35] 169472 128$\times$128$\times$1 7976
5 SimpleNet[38] 166448 32$\times$32$\times$3 9237
6 WideNet[38] 313700 32$\times$32$\times$3 10091
3 7 EfficientNetV2[39] 627220 32$\times$32$\times$3 66468
4 8 MobileNetV2[36] 821164 32$\times$32$\times$3 296318
Figure 14: Example execution plans (workload 2, 4).
Figure 15: Throughput, latency, and power consumption result for four different workloads.
VI-A2 Baselines
@@
IndModel: State-of-the-art model partitioning methods [10, 11, 12, 13, 14] primarily determine the optimal split execution plan based on metric estimations. However, these methods are designed for single-model optimization and do not directly extend to our multi-model scenario. We categorize them under IndModel and adapt them for our use case. Specifically, IndModel selects the best-split execution plan for each pipeline independently, without a holistic view. The final collaborative execution plan is then formed by aggregating these independently selected plans.
@@
IndE2E: IndE2E advances IndModel by incorporating source and target devices alongside model splitting. Each pipeline independently selects the execution plan expected to yield the highest throughput by considering end-to-end (E2E) latency, from sensing to output delivery. However, it does not account for the resource usage of other pipelines.
VI-A3 Performance Metrics
We consider three performance metrics: throughput, latency, and energy. Throughput is defined as the total number of model inferences per second, while latency is the time taken to execute an end-to-end holistic collaboration plan, respectively. For energy, we report the average power consumption for the execution of a holistic collaboration plan. We measured power consumption with Monsoon power monitor [40].
VI-B Overall Performance
Figure 15 shows the result of different methods for the four workloads. In terms of throughput, Synergy consistently outperforms all baselines owing to its holistic decision-making process and adaptive task parallelization. Synergy on average shows 23.0$\times$ higher throughput than the baselines. In Workload 1 and 2 where multiple pipelines run concurrently, IndModel (state-of-the-art model partitioning methods) results in OOR situations. In Workload 3 and 4, Synergy continues to exhibit 1.8$\times$ and 2.2$\times$ higher throughput than the second best (IndE2E) respectively. The results indicate two key points. First, Synergy boosts throughput without OOR failures when multiple pipelines exist (Workload 1 and 2). Second, Synergy supports large models through the collaboration of distributed AI accelerators (Workload 3 and 4).
@@ VI-B Overall Performance
TABLE II: Ablation study of Synergy.
JRC STT PSR ATP Workload 1 Workload 2
TPUT (inf/s) Latency (s) Power (J/s) TPUT (inf/s) Latency (s) Power (J/s)
OOR OOR OOR OOR OOR OOR
0.06 47.63 1.52 2.30 1.31 1.55
0.92 3.25 1.59 15.28 0.20 1.69
2.72 1.10 1.61 15.28 0.20 1.69
4.20 0.86 1.47 29.67 0.10 1.75
VI-B1 Ablation Study
We perform an ablative study to understand the effectiveness of each technical component of Synergy. Table II shows the result. JRC refers to the joint resource consideration among pipelines, and STT refers to device mapping accounting for source and target tasks (§IV-C). PSR means progressive search space reduction (§IV-D), and ATP means adaptive task parallelization (§IV-F). If none of these are applied, it is the same as the state-of-the-art approach (IndModel). If all are checked, it becomes Synergy. Overall, the throughput and latency improve as each of our components is added. Improving IndModel by addressing resource conflicts among pipelines to avoid OOR (JRC) is the same as JointModel in our baselines. Note that merely resolving the resource conflicts is still far from Synergy’s performance. Incorporating other technical components, STT, PSR, and ATP shows 6.9, 7.6, and 14.3$\times$ higher throughput compared to JRC-only, respectively. Similarly, the latency was reduced by 93.0%, 97.3%, and 98.0% compared to JRC-only, owing to the parallelization of the pipelines. We also noted that utilizing all components increased power consumption by around 12.9% compared to JRC-only, due to the increased use of computation units. These findings are aligned with the earlier analysis of the overall performance.
VI-C In-Depth Analysis
VI-C1 Runtime Environment Changes
@@
Figure 16: Impact of runtime environment changes: (a) number of devices and (b) number of pipelines.
Number of devices: In this experiment, we explore Synergy’s scalability to the change of the number of MAX78000 devices. We increase the number of devices from two to five while running the same set of four pipelines with ConvNet5, KWS, SimpleNet, and ResSimpleNet. Figure 16 (a) shows the result. Interestingly, Synergy significantly outperforms the baselines as the number of devices increases. This is due to Synergy’s strategic consideration of accelerator assignments, effectively minimizing communication overhead between source and target devices. Conversely, except for IndE2E, the throughput gains for all other baselines are not notable, even with additional devices. Another interesting observation is that using more devices does not always lead to higher throughput. In the case of Synergy, the throughput saturates once the number of devices reaches 4. This is because, beyond optimal distribution, further splitting of models fails to contribute to additional throughput gains.
@@
VI-C2 Composing Heterogeneous Accelerators
Figure 17: Effect of accelerator composition.
To examine Synergy’s effectiveness in heterogeneous accelerator resources, we conducted an experiment where one of four MAX78000 devices was substituted with more resource-capable MAX78002. The workload is comprised of three pipelines: ConvNet5, UNet, and EfficientNet. Figure 17 shows the results in two different setups. The inclusion of a higher-resource device generally led to an improvement in throughput. With four MAX78000s, Synergy achieves a total throughput of 0.93, outperforming the second best (JointModel) by a factor of 16.4. This throughput further increases to 3.33 when incorporating one MAX78002. PriMinDev, which allocates all models exclusively to the single MAX78002, results in a significantly low throughput of 0.06. This discrepancy demonstrates why simply offloading model execution to a more powerful device (MAX78002) is not always the best approach. It also underscores the need to account for communication overhead between source and target devices. Additionally, we observed that while IndE2E results in an OOR condition with four MAX78000s, it achieves the second-highest throughput when using the MAX78002. This suggests that IndE2E performs well in resource-rich environments but fails to manage resources carefully in resource-constrained settings.
Figure 18: Effect of source and target mappings.
VI-C3 Source and Target Mapping
We assessed source and target mapping effects in three scenarios: (i) any device can be source or target (Any), (ii) source/target devices are evenly allocated (Distributed), and (iii) the same device is used as both source and target across pipelines (Overlapped). The Distributed setup matches Workload 1 from §VI-B, while the Overlapped and Any scenarios only differ in their source and target device mappings. As shown in Figure 18, the Overlapped scenario has the lowest throughput due to communication bottlenecks, while the Any scenario has the highest throughput by distributing communication costs. The result indicates Synergy’s robustness in execution planning, considering both source and target factors.
TABLE III: Comparison among different objectives.
@@ VI-C4 Different Objectives
While Synergy defaults to maximizing system-wide throughput (TPUT-max), it can also minimize latency (Latency-min) or power consumption (Power-min). Table III demonstrates Synergy’s ability to achieve these different objectives effectively. Each objective excels in its metric, though identical latencies do not ensure the same holistic collaboration plan. Notably, maximizing throughput provides a balanced performance with low latency and near-minimal power use. For example, in workload 1, TPUT-max achieves 22.1$\times$ higher throughput with only 1.2$\times$ more power consumption compared to Power-min. This is because maximizing throughput reduces latency and communication overhead, the primary power consumption source.
Figure 19: Effect of the objective: minimizing power consumption.
We further investigate the effectiveness of Synergy’s orchestration by taking Power-min as an example. Figure 19 shows the average power consumption for Workload 1 and Workload 2 when Power-min is set as an objective. The baselines also select the execution plan, prioritizing the minimization of power consumption. The results show that Synergy executes multiple workloads while minimizing power consumption and avoiding OOR.
VII Related work
TinyML: TinyML represents a research field in machine learning techniques, aiming to bring AI capabilities to the most resource-constrained devices, such as MCUs. These devices typically have tens to hundreds of kilobytes of SRAM, and most of the research efforts in this domain have focused on minimizing model size. Existing studies have explored mainly three techniques: model pruning [41], model quantization [42], and neural architecture search (NAS) [43, 44]. Model pruning [41] reduces model complexity by identifying and eliminating unnecessary parameters that contribute minimally to output accuracy. Model quantization [42] decreases both model size and computational intensity of operations by reducing the precision of the numerical values used to represent a model’s parameters. NAS [43, 44] is an automated method for discovering the best neural network architectures that are tailored to the resource constraints of a specific device, while balancing efficiency and accuracy. In contrast, our work focuses on supporting multiple or large AI models without compromising accuracy, by dynamically composing distributed AI accelerators. However, the combined power of tiny AI accelerators may still be insufficient to support unmodified, off-the-shelf large AI models. We envision that Synergy can benefit from these TinyML techniques to accommodate larger AI models.
@@ VII Related work
Model partitioning: While there are very few attempts for partitioning AI models over multiple MCU-equipped devices, there have been active research efforts for layer-wise model partitioning over resource-limited embedded and mobile devices. In common, these methods [10, 11, 12, 13, 14, 15, 16] allocates few initial layers of a DNN on mobile/embedded device and the latter in edge or cloud server. Intermediate output from the initial layer execution is transmitted to powerful resources such as the cloud or nearby edge device where the subsequent part of the model is executed. They adapt the splitting layer depending on network status and server load. While Synergy also adopts vertical partitioning for distributed inferences, existing model partitioning studies lack the (i) end-to-end perspective of the on-body AI app execution (sensor and interface operations as well as model execution) and (ii) consideration of concurrent execution of multi-tenant models. Our experiments demonstrate that Synergy outperforms these studies by solving the challenges.
@@
DNN workload distribution: An application pipeline often consists of a (conditional) sequence of multiple DNN models. Several distributed systems have been proposed to balance model workloads across distributed devices [45, 46, 47]. Their key difference with model partitioning methods is to treat a model as a primitive execution unit and focus more on scheduling model execution over distributed devices (without model splitting). For example, a face classification task can be executed on a different device following object detection on the initial device, thereby leveraging distributed resources from multiple devices [45, 46]. Synergy shares the same high-level objective of distributing multiple model workloads into multiple devices. However, our target environment is on-body devices with AI accelerators, which brings resource challenges. To overcome the limited resources of these devices, we adopt model partitioning on top of workload distribution and devise a tailored solution for tiny AI accelerators.
@@
Model serving systems: Several platforms, such as TensorFlow serving [48], Sagemaker [49], and Azure ML [50], have been proposed to facilitate model inference serving by offering containerized environments for model execution in diverse devices. Research platforms, such as Velox [51] and Clipper [52], focus on low-latency prediction serving, together with optimizing cloud server performance. However, in wearable computing, the main challenge lies in dynamically composing distributed tiny AI accelerators. Our device-agnostic programming interface addresses this by mapping software logic to physical resources and optimizing performance by considering pipeline interdependencies and resource usage.
Middleware for body sensor networks: Several decades ago, body sensor networks (BSNs) received attention for recognizing human contexts using sensory signals from multiple wearable devices. Early work focused on fusion techniques to effectively and robustly concatenate multiple data streams [53, 54, 55, 56, 57] when different devices introduced several challenges, e.g., time synchronization, missing data, and varying sampling rates. Another focus was handling dynamic device availability [58, 59, 60, 61, 62], i.e., when the set of available devices changes over time. To tackle this problem, many studies have explored the effects of different characteristics on recognition accuracy, e.g., types, compositions, and placements of sensors, and have proposed methods to dynamically select the best sensor based on predefined parameters.
Although these works provide a foundation for sensing and processing in multi-device environments, their consideration of sensing pipelines and target devices has been limited to support on-body AI apps we envision. First, Sensing pipelines in BSNs was simple, so pipeline partitioning across devices was not actively studied. Second, due to insufficient processing capabilities, wearable devices were primarily treated as sensor data streamers, with most processing assumed to be run on a smartphone. However, we argue that on-body sensing pipelines have evolved with the introduction of numerous DNN-based AI models, and wearable devices have begun to feature on-device AI capabilities, with the emergence of tiny AI accelerators. This insight necessitates revisiting existing strategies for on-body sensing and processing in multi-wearable environments. To the best of our knowledge, our work is the first to efficiently support on-body AI applications on wearable devices equipped with tiny AI accelerators.
VIII Conclusion
@@
Future work aims to enhance collaboration among AI accelerators by expanding from layer-wise model splitting to include channel-wise splitting [63, 64, 65, 66, 67, 68], enabling Synergy to handle larger inputs and optimize performance through a combination of both techniques [69]. Additionally, we plan to extend Synergy to support more complex model architectures, such as sensor fusion and conditional chaining of multiple models, to meet real-world application demands. Lastly, we aim to enrich the device-agnostic programming interface by adding features like detailed taxonomy for sensor and target device mapping and supporting customized functions, thereby providing developers with greater flexibility in designing on-body AI apps.